OpenStudy (anonymous):

How many solutions does the equation have? A. one solution B. infinite solutions C. no solution

OpenStudy (anonymous):

What equation?

OpenStudy (anonymous):

3(4X-2=(6X-3

OpenStudy (anonymous):

\[3(4x-2)=6x-3\]?

OpenStudy (anonymous):

3(4X-2=(6X-3) Yes

OpenStudy (aakashsudhakar):

12x - 6 = 6x - 3 2x - 1 = 2x - 1 As x can take on any value to make both sides of the equation valid, this problem has infinite solutions.

OpenStudy (anonymous):

Thanks

OpenStudy (aakashsudhakar):

Best Response please!